# SAMba-UNet: Synergizing SAM2 and Mamba in UNet with Heterogeneous Aggregation for Cardiac MRI Segmentation

### v1

@@ Abstract
To address the challenge of complex pathological feature extraction in automated cardiac MRI segmentation, this study proposes an innovative dual-encoder architecture named SAMba-UNet. The framework achieves cross-modal feature collaborative learning by integrating the vision foundation model SAM2, the state-space model Mamba, and the classical UNet. To mitigate domain discrepancies between medical and natural images, a Dynamic Feature Fusion Refiner is designed, which enhances small lesion feature extraction through multi-scale pooling and a dual-path calibration mechanism across channel and spatial dimensions. Furthermore, a Heterogeneous Omni-Attention Convergence Module (HOACM) is introduced, combining global contextual attention with branch-selective emphasis mechanisms to effectively fuse SAM2’s local positional semantics and Mamba’s long-range dependency modeling capabilities. Experiments on the ACDC cardiac MRI dataset demonstrate that the proposed model achieves a Dice coefficient of 0.9103 and an HD95 boundary error of 1.0859 mm, significantly outperforming existing methods, particularly in boundary localization for complex pathological structures such as right ventricular anomalies. This work provides an efficient and reliable solution for automated cardiac disease diagnosis, and the code will be open-sourced.
[…]
1 Introduction
@@
Cardiovascular diseases remain one of the leading causes of global mortality, with early diagnosis predominantly relying on imaging modalities such as cardiac magnetic resonance imaging (MRI). However, conventional MRI analysis requires manual annotation of cardiac structures (e.g., ventricles, myocardium) by specialized clinicians, a process that is time-consuming and prone to subjective variability. This limitation becomes particularly evident in detecting complex pathologies like ischemic heart failure, hypertrophic cardiomyopathy, and right ventricular anomalies, where human interpretation struggles to achieve high sensitivity and consistency. Driven by transformative advances in artificial intelligence (AI), the medical community is now developing automated algorithmic systems to enhance diagnostic efficiency Sanyaolu (2025), minimize human error, and deliver data-driven decision support for early screening, personalized treatment optimization, and prognostic evaluation of cardiac disorders.
@@
In recent years, deep learning architectures have achieved remarkable breakthroughs in medical image segmentation, with encoder-decoder based convolutional neural networks (CNNs) demonstrating exceptional performance Ronneberger et al. (2015); Zhou et al. (2019). As a milestone architecture in this field, UNet has established its core position in medical image segmentation tasks through its unique symmetric encoder-decoder design and cross-level skip connection mechanism. To enhance feature representation capabilities, subsequent studies have developed various innovative auxiliary modules Huang et al. (2017); He et al. (2016); Woo et al. (2018); Howard et al. (2017); Yu and Koltun (2015); Zhou et al. (2020). These technological advancements have empowered the UNet architecture to demonstrate outstanding clinical value in segmenting medical images across multiple modalities, including CT, MRI, and ultrasound imaging.
@@
Transformer demonstrates significant advantages in modeling long-range dependencies and capturing global context through its attention mechanism Chen et al. (2021); Zhang et al. (2021); Valanarasu et al. (2021); Hatamizadeh et al. (2021). Driven by advanced network architectures Dosovitskiy et al. (2020) and large-scale datasets Kirillov et al. (2023), recent segmentation trends have shifted from task-specific expert models toward general-purpose foundation models that can perform segmentation without extensive task-specific development Moor et al. (2023); He et al. (2024); Khan et al. (2025). SAM Kirillov et al. (2023) and SAM2 Ravi et al. (2024), as newly developed visual foundation models, have shown impressive zero-shot performance across various natural image tasks. However, the substantial domain gap between natural images and MRI scans prevents direct deployment of SAM on medical imaging Huang et al. (2024); He et al. (2023); Roy et al. (2023). To address the challenges of blurred or missing small lesions and fine structures when applying SAM2 to MRI segmentation due to its training on natural images, we propose a Dynamic Feature Fusion Refiner in this study.
@@
Although the Transformer architecture demonstrates remarkable advantages in modeling long-range dependencies, its inherent quadratic computational complexity leads to excessive resource consumption in medical image segmentation tasks Liu et al. (2024a); Huang et al. (2022). In contrast, the Mamba architecture emerges as a promising solution for medical imaging due to its linear computational complexity and powerful capability in capturing long-range dependencies Xing et al. (2024); Yang et al. (2024). In the Hiera architecture employed by SAM2, the window attention mechanism and window-based absolute positional encoding may cause loss of pixel-level spatial positional semantic information, thereby compromising the extraction of segmentation boundary features. To enhance long-range dependency learning while reducing computational resource consumption, we introduce the VMamba architecture to capture global semantic features that complement SAM2’s hierarchical features. Addressing the challenge of adaptive fusion between SAM2 and Mamba encoder features, we innovatively design a Heterogeneous Omni-Attention Convergence Module (HOACM) to effectively integrate heterogeneous semantic features from both architectures. Based on our collection and analysis of existing data, it indicates that the model we have proposed, SAMba-UNet represents the first pioneering framework that successfully synergizes SAM2, Mamba, and UNet architectures.
@@
2 Related Work
2.1 Medical Image Segmentation with SAM
@@
The introduction of the Segment Anything Model (SAM) Kirillov et al. (2023) represents a significant milestone in image segmentation. While SAM demonstrates remarkable zero-shot segmentation capabilities through user prompts (e.g., points, bounding boxes) without task-specific training, its direct application to medical imaging, particularly MRI analysis, faces challenges due to substantial domain discrepancies between natural images and medical imaging modalities. To address this, parameter-efficient fine-tuning (PEFT) Zu et al. (2024) strategies offer an effective solution by updating minimal parameters (typically <5% of total weights) while maintaining most parameters frozen. Building on this concept, Wu et al. (2025) proposes a Medical SAM Adapter (Med-SA) that injects medical domain knowledge through PEFT rather than direct SAM fine-tuning, achieving state-of-the-art performance in medical image segmentation with only 2% parameter updates. Further optimizing for MRI segmentation tasks, we introduce a Dynamic Feature Fusion Refiner module to address the limitations of SAM in capturing small lesions and subtle anatomical structures, which stems from its extensive training on natural image datasets.
2.2 Medical Image Segmentation with Mamba
@@
State space sequence models (SSMs) such as Mamba Gu and Dao (2023) offer a novel approach for efficient global dependency modeling through their linear complexity (O(n)) in long sequence processing. Unlike self-attention mechanisms, SSMs achieve interaction between sequence elements and historical information by compressing hidden states, thereby avoiding quadratic computational overhead. The Mamba-Unet Wang et al. (2024) framework proposes a novel medical image segmentation model by integrating the U-Net Ronneberger et al. (2015) architecture with Mamba’s capabilities. However, its insufficient local feature extraction capability limits effective capture of subtle lesion structures. Subsequent improvements like U-Mamba Ma et al. (2024) and SegMamba Xing et al. (2024) combine Mamba with CNNs for direct pixel-level long-range dependency modeling. Nevertheless, these Mamba-based models compromise the spatial continuity of local neighborhood pixels due to their 1D sequential processing, adversely affecting detail modeling. To address these limitations, we propose leveraging SAM2’s Ravi et al. (2024) capability in capturing window-based absolute positional spatial semantic information to compensate for the detailed modeling deficiencies inherent in Mamba architectures.
2.3 Synergizing Segment Anything Model with Mamba Architecture
@@
SAM-Mamba Dutta et al. (2025) proposes a Mamba-guided Segment Anything Model for efficient polyp segmentation, introducing a Mamba-Prior module as a bridge to connect SAM’s general pre-trained representations with polyp-related subtle cues. LFSamba Liu et al. (2024b) develops a novel multi-focus light field image salient object detection model that reconstructs 3D scenes using single-focus images to capture spatial geometric information. While existing works integrate SAM with Mamba, this study innovatively combines SAM2 with Mamba. SAM2’s MAE He et al. (2022) -pretrained hierarchical Hiera image encoder Ryali et al. (2023) completely removes relative position bias (RPB) during attention computation, adopting window-based absolute position encoding Yu et al. (2024), which may compromise pixel-level spatial positional semantics and adversely affect segmentation performance. To address this, we propose leveraging Mamba’s significant advantage in capturing global semantic information with linear computational complexity, thereby synergistically enhancing SAM2’s capacity to model both global semantics and position-sensitive local features.
3 Method
@@
Figure 1: The architecture of SAMba-UNet.
3.1 Architecture Overview
Dual-stream encoders In the SAM2 encoder branch, we integrate design principles from FE-UNet and Medical SAM Adapter to establish a dual-Adapter architecture for Hiera Block fine-tuning. To bridge the domain gap between SAM2’s natural image pre-training and MRI characteristics, we propose a Dynamic Feature Fusion Refiner that performs domain-adaptive refinement on attention outputs, complemented by parallel MLP-Adapter operations to jointly enhance nonlinear mapping capacity.
@@ 3.1 Architecture Overview
Single VMamba decoder In the decoder design, we maintain architectural consistency with Mamba-UNet by implementing its progressive upsampling framework. This multi-stage feature aggregation mechanism enables hierarchical feature fusion through successive transpose convolutions, ultimately producing high-resolution segmentation masks via parameter-shared prediction heads.
3.2 Dynamic Feature Fusion Refiner
@@
The Dynamic Feature Fusion Refiner takes attention outputs as input features $X\in R^{B\times H\times W\times C}$ It employs dual adaptive pooling operations: adaptive max pooling preserves globally salient semantics while adaptive average pooling suppresses high-frequency noise interference. These pooled features are concatenated along the channel dimension after dimension permutation:
$X_{\mathrm{cat}}=\mathrm{Concat}([\text{AdaptiveAvgPool}2\mathrm{d}(\mathrm{Permute}(X)),\text{AdaptiveMaxPool}2\mathrm{d}(\mathrm{Permute}(X))],\mathrm{dim}=1)$ (1)
@@
The parameters $W_{1}^{(c)}\in\mathbb{R}^{h\times 2C}$ and $W_{2}^{(c)}\in\mathbb{R}^{C\times h}$ in the formula represent learnable weight matrices, where $h=\lfloor C\cdot r\rfloor$ denotes the bottleneck dimension with compression ratio $r\in[0,1]$ , and $\sigma$ serves as the channel gating activation. A cascaded convolution path is constructed to augment local receptive fields: the downsampling convolution compresses spatial dimensions, while the transposed convolution restores resolution, with ReLU non-linearities enhancing local feature extraction:
@@
3.3 Heterogeneous Omni-Attention Convergence Module
Figure 3: The architecture of HOACM.
@@ 3.3 Heterogeneous Omni-Attention Convergence Module
Bifurcated Selective Emphasis Attention(BSEA) The Bifurcated Selective Emphasis Attention (BSEA) employs a dual-path pooling-attention architecture: 1) A spatial average pooling pathway preserves VMamba’s global context modeling strength by capturing cross-region long-range dependencies, while 2) a local saliency pooling pathway applies spatial constraints to prevent attention over-focusing, thereby alleviating local-global semantic misalignment.
@@
[…]
@@
Omniscient Contextual Attention To address the SAM2 encoder’s limitations in modeling global pixel-level positional semantics, we propose the Omniscient Contextual Attention (OCA) mechanism. This module enhances feature representation through a global contextual awareness attention architecture that establishes long-range cross-region dependencies and enables multi-scale semantic integration:
@@
A gated spatial attention (GSA) mechanism then models long-range spatial dependencies using 7×7 convolutional kernels:
@@
4 Experiments and Results
Automated Cardiac Diagnosis Challenge The ACDC dataset originates from the 2017 MICCAI challenge of the same name, comprising cardiac MRI short-axis sequences from 100 patients collected by multiple French clinical centers, including the University Hospital of Dijon. This dataset encompasses five cardiac pathologies and normal cases: dilated cardiomyopathy (DCM) characterized by left ventricular enlargement, hypertrophic cardiomyopathy (HCM) marked by abnormal thickening of left ventricular myocardium, myocardial infarction (MINF) presenting left ventricular myocardial scarring, right ventricular dysfunction (RV-abnormal) manifesting structural/contractile abnormalities in the right ventricle, and normal cardiac anatomy. Pathological features primarily localize in the left ventricle (DCM, HCM, MINF) or right ventricle (RV-abnormal), requiring comprehensive multi-level cardiac segmentation and functional analysis through MRI short-axis views for accurate diagnosis.
4.1 Implementation Details
The experimental setup was established on an Ubuntu 23.10 operating system, utilizing Python 3.12.0 with PyTorch 1.10 deep learning framework accelerated by CUDA 12.1. The hardware configuration consisted of an NVIDIA A800-SXM4-80GB GPU and an Intel Xeon Platinum 8462Y+ CPU. We employed the preprocessed ACDC dataset for 2D medical image segmentation tasks. The SAMba-UNet model underwent 10,000 training iterations with a batch size of 12. Stochastic Gradient Descent (SGD) optimizer [2] was implemented with an initial learning rate of 0.01, momentum of 0.9, and weight decay coefficient of 0.0001. Model evaluation on the validation set was performed every 200 iterations, accompanied by checkpoint saving of optimal parameters.
4.2 Evaluation Metrics

### v2

# SAMba-UNet: SAM2–Mamba UNet for Cardiac MRI in Medical Robotic Perception

@@ Abstract
To address complex pathological feature extraction in automated cardiac MRI segmentation, we propose SAMba-UNet, a novel dual-encoder architecture that synergistically combines the vision foundation model SAM2, the linear-complexity state-space model Mamba, and the classical UNet to achieve cross-modal collaborative feature learning; to overcome domain shifts between natural images and medical scans, we introduce a Dynamic Feature Fusion Refiner that employs multi-scale pooling and channel–spatial dual-path calibration to strengthen small-lesion and fine-structure representation, and we design a Heterogeneous Omni-Attention Convergence Module (HOACM) that fuses SAM2’s local positional semantics with Mamba’s long-range dependency modeling via global contextual attention and branch-selective emphasis, yielding substantial gains in both global consistency and boundary precision—on the ACDC cardiac MRI benchmark, SAMba-UNet attains a Dice of 0.9103 and HD95 of 1.0859 mm, notably improving boundary localization for challenging structures like the right ventricle, and its robust, high-fidelity segmentation maps are directly applicable as a perception module within intelligent medical and surgical robotic systems to support preoperative planning, intraoperative navigation, and postoperative complication screening; the code will be open-sourced to facilitate clinical translation and further validation.
I Introduction
@@
Cardiovascular diseases remain one of the leading causes of global mortality, with early diagnosis predominantly relying on imaging modalities such as cardiac magnetic resonance imaging (MRI). However, conventional MRI analysis requires manual annotation of cardiac structures (e.g., ventricles, myocardium) by specialized clinicians, a process that is time-consuming and prone to subjective variability. This limitation becomes particularly evident in detecting complex pathologies like ischemic heart failure, hypertrophic cardiomyopathy, and right ventricular anomalies, where human interpretation struggles to achieve high sensitivity and consistency. Driven by transformative advances in artificial intelligence (AI), the medical community is now developing automated algorithmic systems to enhance diagnostic efficiency [1], minimize human error, and deliver data-driven decision support for early screening, personalized treatment optimization, and prognostic evaluation of cardiac disorders. In addition, these image-driven AI systems are increasingly being integrated with surgical/medical robotics to support image‑guided minimally invasive procedures, intraoperative navigation, and automated assistance in clinical decision making, thereby enabling a longitudinal closed‑loop from early screening to interventional treatment.
@@
In recent years, deep learning architectures have achieved remarkable breakthroughs in medical image segmentation, with encoder-decoder based convolutional neural networks (CNNs) demonstrating exceptional performance [2, 3]. As a milestone architecture in this field, UNet has established its core position in medical image segmentation tasks through its unique symmetric encoder-decoder design and cross-level skip connection mechanism. To enhance feature representation capabilities, subsequent studies have developed various innovative auxiliary modules [4, 5, 6, 7, 8, 9]. These technological advancements have empowered the UNet architecture to demonstrate outstanding clinical value in segmenting medical images across multiple modalities, including CT, MRI, and ultrasound imaging.
@@
Transformer demonstrates significant advantages in modeling long-range dependencies and capturing global context through its attention mechanism [10, 11, 12, 13]. Driven by advanced network architectures [14] and large-scale datasets [15], recent segmentation trends have shifted from task-specific expert models toward general-purpose foundation models that can perform segmentation without extensive task-specific development [16, 17, 18]. SAM [15] and SAM2 [19], as newly developed visual foundation models, have shown impressive zero-shot performance across various natural image tasks. However, the substantial domain gap between natural images and MRI scans prevents the direct deployment of SAM on medical imaging [20, 21, 22, 23]. To address the challenges of blurred or missing small lesions and fine structures when applying SAM2 to MRI segmentation due to its training on natural images, we propose a Dynamic Feature Fusion Refiner in this study.
Figure 1: Workflow of the proposed medical robotic control system.
@@
Although the Transformer architecture demonstrates remarkable advantages in modeling long-range dependencies, its inherent quadratic computational complexity leads to excessive resource consumption in medical image segmentation tasks [24, 25]. In contrast, the Mamba architecture emerges as a promising solution for medical imaging due to its linear computational complexity and powerful capability in capturing long-range dependencies [26, 27]. In the Hiera architecture employed by SAM2, the window attention mechanism and window-based absolute positional encoding may cause the loss of pixel-level spatial positional semantic information, thereby compromising the extraction of segmentation boundary features. To enhance long-range dependency learning while reducing computational resource consumption, we introduce the VMamba architecture to capture global semantic features that complement SAM2’s hierarchical features. Addressing the challenge of adaptive fusion between SAM2 and Mamba encoder features, we innovatively design a Heterogeneous Omni-Attention Convergence Module (HOACM) to effectively integrate heterogeneous semantic features from both architectures. Based on our collection and analysis of existing data, it indicates that the model we have proposed, SAMba-UNet, represents the first pioneering framework that successfully synergizes SAM2, Mamba, and UNet architectures.
Figure 1 illustrates the application workflow of an intelligent robotic control system in the medical field. Key components include: a patient undergoing CT scanning with robotic assistance on the left; the intermediate image processing stages featuring VMamba Branch for global semantic feature extraction and SAM2 Branch for spatial positional feature extraction; and the three-phase intelligent robot control system (Preoperative, Intraoperative, Postoperative) on the right.
@@
II Related Work
II-A Medical Image Segmentation with SAM
@@
The introduction of the Segment Anything Model (SAM) [15] represents a significant milestone in image segmentation. While SAM demonstrates remarkable zero-shot segmentation capabilities through user prompts (e.g., points, bounding boxes) without task-specific training, its direct application to medical imaging, particularly MRI analysis, faces challenges due to substantial domain discrepancies between natural images and medical imaging modalities. To address this, parameter-efficient fine-tuning (PEFT) [28] strategies offer an effective solution by updating minimal parameters (typically <<5% of total weights) while maintaining most parameters frozen. Building on this concept, [29] proposes a Medical SAM Adapter (Med-SA) that injects medical domain knowledge through PEFT rather than direct SAM fine-tuning, achieving state-of-the-art performance in medical image segmentation with only 2% parameter updates. Further optimizing MRI segmentation, we introduce a Dynamic Feature Fusion Refiner to mitigate SAM’s limitations in capturing small lesions and subtle anatomical structures, and we add a Mamba branch to obtain long-range context with linear complexity—enabling efficient modeling of cross-region dependencies in cardiac MRI.
II-B Medical Image Segmentation with Mamba
Figure 2: The architecture of SAMba-UNet.
@@ II-B Medical Image Segmentation with Mamba
State space sequence models (SSMs) such as Mamba [30] offer a novel approach for efficient global dependency modeling through their linear complexity (O(n)) in long sequence processing. Unlike self-attention mechanisms, SSMs achieve interaction between sequence elements and historical information by compressing hidden states, thereby avoiding quadratic computational overhead. The Mamba-Unet [31] framework proposes a novel medical image segmentation model by integrating the U-Net [2] architecture with Mamba’s capabilities. However, its insufficient local feature extraction capability limits effective capture of subtle lesion structures. Subsequent improvements like U-Mamba [32] and SegMamba [26] combine Mamba with CNNs for direct pixel-level long-range dependency modeling. Nevertheless, these Mamba-based models compromise the spatial continuity of local neighborhood pixels due to their 1D sequential processing, adversely affecting detail modeling. To address these limitations, we propose leveraging SAM2’s [19] capability in capturing window-based absolute positional spatial semantic information to compensate for the detailed modeling deficiencies inherent in Mamba architectures.
II-C Synergizing SAM with Mamba Architecture
@@
SAM-Mamba [33] proposes a Mamba-guided Segment Anything Model for efficient polyp segmentation, introducing a Mamba-Prior module as a bridge to connect SAM’s general pre-trained representations with polyp-related subtle cues. LFSamba [34] develops a novel multi-focus light field image salient object detection model that reconstructs 3D scenes using single-focus images to capture spatial geometric information. While existing works integrate SAM with Mamba, this study innovatively combines SAM2 with Mamba. SAM2’s MAE [35] -pretrained hierarchical Hiera image encoder [36] completely removes relative position bias (RPB) during attention computation, adopting window-based absolute position encoding [37], which may compromise pixel-level spatial positional semantics and adversely affect segmentation performance. To address this, we propose leveraging Mamba’s significant advantage in capturing global semantic information with linear computational complexity, thereby synergistically enhancing SAM2’s capacity to model both global semantics and position-sensitive local features.
III The Propose Method
@@
III-A Architecture Overview
Dual-stream Encoders. In the SAM2 encoder branch, we integrate design principles from FE-UNet and Medical SAM Adapter to establish a dual-Adapter architecture for Hiera Block fine-tuning. To bridge the domain gap between SAM2’s natural image pre-training and MRI characteristics, we propose a Dynamic Feature Fusion Refiner that performs domain-adaptive refinement on attention outputs, complemented by parallel MLP-Adapter operations to jointly enhance nonlinear mapping capacity.
@@ III-A Architecture Overview
Single VMamba Decoder. In the decoder design, we maintain architectural consistency with Mamba-UNet by implementing its progressive upsampling framework. This multi-stage feature aggregation mechanism enables hierarchical feature fusion through successive transpose convolutions, ultimately producing high-resolution segmentation masks via parameter-shared prediction heads.
III-B Dynamic Feature Fusion Refiner
@@
The Dynamic Feature Fusion Refiner takes attention outputs as input features $X\in R^{B\times H\times W\times C}$. It employs dual adaptive pooling operations: adaptive max pooling preserves globally salient semantics while adaptive average pooling suppresses high-frequency noise interference. These pooled features are concatenated along the channel dimension after dimension permutation:
$\begin{split}X_{\mathrm{cat}}=\mathrm{Concat}([&\text{AdaptiveAvgPool}2\mathrm{d}(\mathrm{Permute}(X)),\\ &\text{AdaptiveMaxPool}2\mathrm{d}(\mathrm{Permute}(X))],\mathrm{dim}=1)\end{split}$ (1)
@@
The parameters $W_{1}^{(c)}\in\mathbb{R}^{h\times 2C}$ and $W_{2}^{(c)}\in\mathbb{R}^{C\times h}$ in the formula represent learnable weight matrices, where $h=\lfloor C\cdot r\rfloor$ denotes the bottleneck dimension with a compression ratio $r\in[0,1]$ , and $\sigma$ serves as the channel gating activation. A cascaded convolution path is constructed to augment local receptive fields: the downsampling convolution compresses spatial dimensions, while the transposed convolution restores resolution, with ReLU non-linearities enhancing local feature extraction:
@@
III-C Heterogeneous Omni-Attention Convergence Module
Figure 4: The architecture of eterogeneous Omni-Attention Convergence Module (HOACM).
@@ III-C Heterogeneous Omni-Attention Convergence Module
Bifurcated Selective Emphasis Attention (BSEA). The Bifurcated Selective Emphasis Attention (BSEA) employs a dual-path pooling-attention architecture: 1) A spatial average pooling pathway preserves VMamba’s global context modeling strength by capturing cross-region long-range dependencies, while 2) a local saliency pooling pathway applies spatial constraints to prevent attention over-focusing, thereby alleviating local-global semantic misalignment.
@@
$Q$ $=\text{Flatten}(\text{Conv2D}_{1\times 1}(X_{avg/max};W_{q}))$ (7)
$K$ $=\text{Flatten}(\text{Conv2D}_{1\times 1}(X_{avg/max};W_{k}))$ (8)
$V$ $=\text{Flatten}(\text{Conv2D}_{1\times 1}(X_{avg/max};W_{v}))$ (9)
@@
Omniscient Contextual Attention. To address the SAM2 encoder’s limitations in modeling global pixel-level positional semantics, we propose the Omniscient Contextual Attention (OCA) mechanism. This module enhances feature representation through a global contextual awareness attention architecture that establishes long-range cross-region dependencies and enables multi-scale semantic integration:
@@
A gated spatial attention (GSA) mechanism then models long-range spatial dependencies using $\times7$ convolutional kernels:
@@
IV Experiments
IV-A Automated Cardiac Diagnosis Challenge
The ACDC dataset originates from the 2017 MICCAI challenge of the same name, comprising cardiac MRI short-axis sequences from 100 patients collected by multiple French clinical centers, including the University Hospital of Dijon. This dataset encompasses five cardiac pathologies and normal cases: dilated cardiomyopathy (DCM) characterized by left ventricular enlargement, hypertrophic cardiomyopathy (HCM) marked by abnormal thickening of the left ventricular myocardium, myocardial infarction (MINF) presenting left ventricular myocardial scarring, right ventricular dysfunction (RV-abnormal) manifesting structural/contractile abnormalities in the right ventricle, and normal cardiac anatomy. Pathological features primarily localize in the left ventricle (DCM, HCM, MINF) or right ventricle (RV-abnormal), requiring comprehensive multi-level cardiac segmentation and functional analysis through MRI short-axis views for accurate diagnosis.
IV-B Implementation Details
The experimental setup was established on an Ubuntu 23.10 operating system, utilizing Python 3.12.0 with the PyTorch 1.10 deep learning framework accelerated by CUDA 12.1. The hardware configuration consisted of an NVIDIA A800-SXM4-80GB GPU and an Intel Xeon Platinum 8462Y+ CPU. We employed the preprocessed ACDC dataset for 2D medical image segmentation tasks. The SAMba-UNet model underwent 10,000 training iterations with a batch size of 12. Stochastic Gradient Descent (SGD) optimizer [2] was implemented with an initial learning rate of 0.01, a momentum of 0.9, and a weight decay coefficient of 0.0001. Model evaluation on the validation set was performed every 200 iterations, accompanied by checkpoint saving of optimal parameters.
IV-C Evaluation Metrics